# Layer-2 Arbitrage: An Empirical Analysis of Swap Dynamics and Price Disparities on Rollups

### v1

@@ Abstract
This paper explores the dynamics of Decentralized Finance (DeFi) within the Layer-2 ecosystem, focusing on Automated Market Makers (AMM) and arbitrage on Ethereum rollups. We observe significant shifts in trading activity from Ethereum to rollups, with swaps on rollups happening 2–3 times more often, though, with lower trade volume. By examining the price differences between AMMs and centralized exchanges, we discover over 0.5 million unexploited arbitrage opportunities on rollups. Remarkably, we observe that these opportunities last, on average, 10 to 20 blocks, requiring adjustments to the LVR metrics to avoid double-counting arbitrage. Our results show that arbitrage in Arbitrum, Base, and Optimism pools ranges from 0.03% to 0.05% of trading volume, while in zkSync Era it oscillates around 0.25%, with the LVR metric overestimating arbitrage by a factor of five. Rollups offer not only lower gas fees, but also provide faster block production, leading to significant differences compared to the trading and arbitrage dynamics of Ethereum.
Keywords: Arbitrage, Automated Market Maker, Rollup, MEV
@@ 1 Introduction
Ethereum has undergone a series of substantial enhancements aimed at increasing its functionality, addressing scalability challenges, and improving user interactions [2]. A pivotal upgrade was the transition from Proof of Work (PoW) to Proof of Stake (PoS), referred to as The Merge, which occurred throughout 2022 and 2023. After that, the most significant milestone was the Dencun upgrade, implemented on March 13th, 2024. It introduced Proto-Danksharding, specifically designed for rollups: Layer-2 (L2) scaling solutions. This upgrade marked the initial phase of Ethereum’s rollup-centric roadmap and introduced blobs, temporary storage for rollup transaction data, streamlining verification, and reducing data storage costs [2]. Consequently, the gas fees on rollups have been significantly reduced, as shown in Figure 1.
Figure 1: Median transaction fees, measured in Gwei, for selected Ethereum rollups, pre- and post-Dencun Upgrade on March 13th, 2024.
The reduction in gas fees has influenced the use of Decentralized Finance (DeFi) on rollups. In recent years, the Total Value Locked (TVL) and the trading volume on rollups have exhibited a notable increase, with certain rollups now surpassing Ethereum in daily transaction numbers [1, 3]. Furthermore, the dynamics of L2 DeFi is not solely altered by reduced gas fees. The second important factor is the faster block production, typically less than 2 seconds in contrast to Ethereum’s 12 seconds [28]. The combination of lower gas fees and faster block production enables arbitrageurs to capitalize on smaller price discrepancies between trading venues more quickly. Although arbitrage transactions are performed predominantly on Ethereum through Maximal Extractable Value (MEV) Boost Auctions [16], an equivalent MEV boost mechanism is not present in rollups. The absence of MEV boost auctions on rollups can be explained by the fact that most rollups rely on centralized sequencers to aggregate transactions into L1 batches.
@@
The centralized sequencers and the associated censorship risk remain a concern within the L2 DeFi ecosystem. Additional concerns encompass transaction finality, the time period until an L2 transaction becomes irrevocable from the rollup (soft finality) or the underlying layer-1 blockchain (hard finality) [31]. Rollups can be categorized into optimistic and Zero-Knowledge (ZK) rollups based on transaction validity and proof mechanisms [11]. Despite the seven-day hard finality period, optimistic rollups have gained user adoption more rapidly due to better Ethereum Virtual Machine (EVM) compatibility. However, ZK-rollups offer higher transaction compression, throughput, and faster hard finality, ranging from minutes to hours. The aforementioned factors — reduced gas fees, absence of MEV auctions, varying finality speeds, and accelerated block production — contribute to distinct DeFi dynamics on rollups, which warrant further investigation.
@@
With the emergence of decentralized rollup sequencers [26], there arises a potential risk that adverse practices associated with MEV on Ethereum [27] transfer to the L2 DeFi ecosystem. To address this concern, a time-boost mechanism within sequencers has been proposed [22]. This mechanism is specifically designed to allow back-running MEVs while preventing front-running activities. Front-running MEV can facilitate various forms of adversarial actions, such as sandwich attacks on traders [19] and Just-In-Time (JIT) liquidity attacks on Liquidity Providers (LPs) within AMM-DEXs [32]. In contrast, back-running MEV enables immediate arbitrage at DEXs or liquidations at lending protocols. Faster arbitrage execution benefits traders, as it eliminates price discrepancies between trading venues. Additionally, MEV time-boost auctions can generate new revenue for sequencer operators.
This paper contributes to the research on L2 DeFi by empirically examining the differential swap dynamics on rollups and Ethereum, considering the factors mentioned above, such as lower gas fees, faster block production, and the absence of MEV auctions. Furthermore, we quantify the potential arbitrage opportunities across rollups and between rollups and CEXs that can be the basis for the revenue estimation from the time-boost MEV auctions in the rollup sequencers.
Related Work.
@@
There exists a substantial body of research on DeFi, particularly focused on AMM-DEXs. In particular, the systematization of knowledge (SoK) by Xu et al. [30] and empirical investigations of Uniswap (v3) on the Ethereum blockchain [18, 9, 17, 21, 12] are prominent. Most of these studies evaluate the profitability of LPs, specifically addressing impermanent loss. The analysis by Adams et al. [4] delves into the costs and dynamics of swaps on Uniswap (v3) on Ethereum. Our research extends this by breaking down swap fees on rollups and assessing the influence of blobs and block time on swap dynamics. Our findings are consistent with the simulated decomposition of swap fees in rollups presented in Gogol et al. [14].
@@
The loss versus rebalancing metric (LVR), which compares arbitrageur profits with those of LPs, was introduced by Milionis et al. [24, 25] and empirically measured by Fritsch et al. [10]. LVR evaluates the LP portfolio against a rebalancing portfolio that perfectly hedges the market risk of the LP position. Our investigation quantifies the absolute profit of arbitrageurs, also known as the net LVR (per block), incorporating significant adjustments. First, LVR assumes that the price discrepancies between AMMs and target exchanges, such as CEXs, disappear after each block. Our findings confirm this assumption for Uniswap (v3) on Ethereum; nonetheless, the decay period of price discrepancies on rollups extends over multiple blocks. Our methodology adjusts the LVR to this scenario and prevents double counting of the same arbitrage opportunities. Second, LVR and the rebalancing portfolio assess the LP’s loss relative to arbitrageurs, thereby containing impermanent loss. Our metrics measure the absolute profit of the arbitrageur. Practically, arbitrageurs do not replicate the rebalancing portfolio but instead capitalize on arbitrage opportunities [16].
@@
The persistence of price disparities on AMMs within zkSync Era, a ZK rollup, was initially reported by Gogol et al. [13]. Our paper builds on that research by providing: i) a comprehensive analysis of rollups (Arbitrum, Base, Optimism, in addition to zkSync Era) alongside Ethereum, ii) enhanced precision in market price accuracy and per-block analysis, (improving from a 1-minute to a 1-second resolution), and iii) an in-depth examination of Uniswap (v3), focusing on the most actively traded pools.
@@
Heimbach et al. [16] focus on the estimation of arbitrage executed through MEV auctions on Ethereum, specifically addressing non-atomic arbitrage opportunities between AMMs and CEXs. Our research extends this investigation by examining rollups, where MEV auctions have not yet been implemented. Similarly, we establish the minimum threshold for arbitrage opportunities and derive the arbitrage value using AMM formulas. However, this work seeks to quantify the unexploited arbitrage, which occasionally persisted for several minutes, in contrast to the arbitrage realized via MEV on Ethereum. Other studies on arbitrage [7, 20, 8] focus primarily on AMM-DEX on Ethereum and/or cyclic arbitrage opportunities.
@@
Torres et al. [29] is the first effort to quantify MEV in roll-ups. Despite the absence of MEV auctions, this research examines the heuristic for back-running arbitrage MEV, specifically focusing on swap events occurring within the same transaction in the block, within the context of cylindrical arbitrage. Our objective is to assess the potential for non-atomic arbitrage, involving CEXs, or cross-rollup price discrepancies.
@@
Mamageishvili et al.  [22] introduce a time boost mechanism within rollup sequencers. This mechanism facilitates back-running MEV while preventing front-running MEV. In subsequent work [23], a shared sequencer was proposed, enabling cross-rollup MEV auctions, such as those for arbitrage exploitation. However, these studies lack empirical quantification of arbitrage opportunities within rollups, a gap that this work seeks to address.
Contributions.
This paper advances DeFi research within rollups. We conducted an empirical investigation into the distinct swap dynamics in optimistic and ZK rollups, as compared to Ethereum, taking into account specific rollup characteristics, such as lower gas fees, faster block production, and the absence of MEV auctions. Subsequently, we analyze the arbitrage opportunities on rollups that can be exploited by arbitrageurs, thereby providing new revenue streams for sequencers offering back-running MEV auctions. Our principal contributions are as follows.
  • Examination of Arbitrage and Swapping Costs on Rollups: We conduct a comparative study of the costs of swapping and capturing arbitrage on rollups compared to Ethereum. Our findings indicate that, unlike Ethereum, transactions on rollups occur 2–3 times more frequently but are smaller in size. Furthermore, swaps on rollups are generally not impacted by front- or back-running block slippage typically caused by MEV. Block times impact rollup transactions; Ethereum averages one swap per block, while rollups have fewer swaps per block despite higher total swaps. Specifically, swaps occur every third block on Base, fifth on Optimism, and tenth on Arbitrum.
  • Decay of Price Disparities: We show that price differences last for 10–20 blocks in rollups, necessitating modifications in the empirical assessment of LVR metrics. We present a framework for empirically estimating arbitrage in cases where the decay period of price differences spans multiple blocks, capturing only the maximum arbitrage value (MAV) during prolonged price disparities. The empirical MAV method, which filters out repeated arbitrage, indicates that LVR metrics have overestimated arbitrage by a factor of five.
  • Estimation of Arbitrage on Rollups: By examining the price differences between Uniswap (v3) and Binance, we discovered over 0.5 million unexploited arbitrage opportunities on rollups. The MAV metric on Arbitrum, Base, and Optimism pools varies from 0.03% to 0.05% of trading volume, whereas in zkSync Era, it is approximately 0.25%, highlighting significant unresolved price differences.
Paper Organization.
Section 2 offers foundational knowledge on rollups and AMMs. Section 3 details the dataset utilized in our analysis, while Section 4 breaks down swap costs on Ethereum rollups, both pre- and post-Decun (blobs). Section 5 sets the theoretical foundation for arbitrage in AMMs and its estimation framework in rollups. Section 6 presents the results of our empirical investigations. Section 7 examines the significance of arbitrage in rollups and proposes future research directions. Lastly, Section 8 summarizes our findings as well as some final remarks.
@@ 2 Background
This section introduces rollups, a new mechanism of layer-2 blockchain scaling, and formalizes the blockchain terminology used throughout the manuscript. Furthermore, it presents the basics of Decentralized Exchanges (DEXs) and Automated Market Makers (AMMs).
2.1 Layer-2 Blockchains
@@
There are two main approaches — Layer-1 (L1) and Layer-2 (L2) scaling — to tackle blockchain scalability challenges and the associated increases in gas fees. L1 scaling involves creating a new blockchain with its own consensus mechanisms and a unique physical infrastructure to support the network. In contrast, L2 scaling uses a different method: it carries out computations off the main blockchain while recording their results on the underlying chain. The main L2 scaling solutions include state (payment) channels, plasma, validium, and rollups. While plasma and state channels aim to move both data and computation off-chain, rollups and validium are non-custodial: they shift computation off-chain but keep compressed transaction data (rollups) or validity proofs (validiums) on the underlying L1 chain. As a result, rollups and validium require fewer trust assumptions compared to state channels, plasma chains, or L1 networks, with data security and integrity being ensured by the security of the underlying chain.
Rollup.
@@
A rollup functions as a virtual blockchain, producing transaction blocks and recording them on the main chain as compressed batches. However, it operates with the risk of adversarial actions such as stopping block production, generating invalid blocks, withholding data, or engaging in other malicious activities. To ensure the correctness of the transaction state in these untrusted environments, rollups use one of two approaches: the optimistic approach and zero-knowledge (ZK) proofs. Figure 2 illustrates the rollup architecture, highlighting its principal components: sequencers and verifiers. Sequencers consolidate transactions to be rolled up onto the L1 chain, thereby achieving gas fee reductions. Verifiers, which are smart contracts on L1, validate the transaction block created by the sequencer, thus ensuring its correctness. Furthermore, ZK-rollups utilize provers responsible for generating ZK validity proofs.
Figure 2: Architecture of a rollup
Soft and Hard Finality.
@@
Transaction finality refers to the point at which the state of a transaction in the blockchain ledger becomes permanent. Within the context of rollups, there are two types of transaction finality: hard finality, which is the permanent state of a transaction on the L1 chain, and soft finality, which is the irreversible state of a transaction on the L2 chain. [31]
Dencun Upgrade and Blobs.

### v2

# Cross-Rollup MEV:

@@ Abstract
This study quantifies the potential non-atomic MEV on Layer-2 (L2) blockchains by measuring the arbitrage opportunities between cross-rollup and DEX-CEX. Over recent years, we observe a shift in trading activities from Ethereum to rollups, with swaps on rollups occurring 2–3 times more frequently, albeit with lower trade volumes. By analyzing the costs of swap on L2s and price discrepancies cross-rollup and DEX-CEX, we identify more than $500000$ unexplored arbitrage opportunities. In particular, we find that these opportunities persist, on average, for 10 to 20 blocks, necessitating the modification of the Loss Versus Rebalancing (LVR) metric to prevent double-counting. Our findings indicate that the arbitrage opportunities in Arbitrum, Base, and Optimism range between 0.03% and 0.05% of the trading volume, while in the ZKsync Era it fluctuates around 0.25%.
@@ 1 Introduction
Decentralized Finance (DeFi) is transitioning towards rollups, the Layer-2 (L2) scaling solutions for Ethereum. The primary reason is that rollups provide equivalent security guarantees as Ethereum while significantly reducing gas fees [41]. Rollups offload complex computations outside the Ethereum network, storing only compressed batches of transactions on-chain. In March 2023, Ethereum underwent the Dencun upgrade [13], the most substantial change since the Merge, which further lowered gas fees on rollups, as shown in Figure 1. The Dencun upgrade introduced Proto-Danksharding and blobs — temporary data storage specifically designed to optimize rollups performance [14].
[…]
The trading volume on Decentralized Exchanges (DEXs) within L2s is increasing, with certain rollups surpassing Ethereum’s daily swap counts [11, 30]. Then, DEXs within L2s present new opportunities for arbitrageurs aiming to capitalize on price disparities across rollups or between rollups and Centralized Exchange (CEX). However, while arbitrage on Ethereum is currently facilitated through Maximal Extractable Value (MEV) boost auctions [24], such mechanisms are yet unavailable in rollups. Most L2s rely on centralized sequencers to produce blocks and aggregate them into Layer-1 batches, employing the first-come-first-served rule. In addition, block production on L2s usually takes less than 2 seconds, as opposed to Ethereum’s 12-second block time [41]. The combination of reduced gas costs and faster block production allows arbitrageurs to exploit minor price differences at a faster pace.
Arbitrage can be classified into two types: i) atomic arbitrage within DEXs on the same blockchain (cycling arbitrage [43]) or ii) non-atomic arbitrage involving CEX or DEX on a different blockchain [24]. Currently, 71% of the MEV transaction volume on Ethereum is related to arbitrage, and in excess of 60% of that arbitrage volume comes from nonatomic transactions [12, 16, 24]. Cycling arbitrage on L2 can be executed as atomic transactions [43], but non-atomic arbitrage faces a risk of state (price) drift. With the significantly reduced gas fees following the Dencun upgrade (see Figure 1), non-atomic arbitrage can be exploited by posting a large amount of prospective arbitrage transactions that are reverted if unsuccessful. The number of revered transactions on L2s increased from 2% to 6% after the Dencun upgrade [15]. Another strategy used by L2 arbitrageurs is the physical location of centralized sequencers, a service sometimes provided by DeFi solvers [9]. Transactions originating from the same data center as the centralized sequencer are likely to be executed first due to shorter network delays [10, 32, 44]. However, risks associated with centralized sequencers related to censorship remain a concern within L2 DeFi. Additional risks include the transaction finality, the time period until an L2 transaction becomes irrevocable from the rollup (soft finality) or the underlying L1 blockchain (hard finality) [47].
@@
Rollups can be categorized into Optimistic and Zero-Knowledge (ZK) rollups based on transaction validity and proof mechanisms [19]. Despite the seven-day hard finality period, optimistic rollups have gained user adoption more rapidly due to better Ethereum Virtual Machine (EVM) compatibility. However, ZK-rollups offer higher transaction compression, throughput, and faster hard finality, ranging from minutes to hours. The aforementioned factors — reduced gas fees, absence of MEV auctions, varying finality times, and faster block production — contribute to distinct arbitrage dynamics on rollups, which requires further investigation.
@@
With the emergence of decentralized sequencers [38], the adverse practices associated with front-running MEV on Ethereum [10, 40, 42] transfer to L2 DeFi. To address this concern, a time-boost mechanism within sequencers has been proposed [34]. This mechanism is specifically designed to allow only back-running MEVs while preventing front-running. Front-running MEV can enable various forms of adversarial actions, such as sandwich attacks on traders [28, 36] and Just-In-Time (JIT) liquidity attacks on Liquidity Providers (LPs) [48]. In contrast, back-running MEV enables arbitrage in DEXs and liquidations at lending protocols. Faster arbitrage execution benefits traders, as it eliminates price discrepancies between trading venues. Additionally, MEV time-boost auctions can generate new revenue for sequencer operators.
1.1 Contributions
This work contributes to the research on DeFi and MEV on L2s by quantifying the potential of non-atomic arbitrage on rollups. We conduct an empirical investigation into price disparities cross-rollups and measure the decay time. Our findings advance ongoing research on the applications of shared sequencer and time-boost (MEV) auctions. Our key contributions are as follows.
L2 Non-atomic Arbitrage and Its Decay Over Time. By examining the price differences between Binance and Uniswap (v3) WETH-USDC pools on L2s, we discover over 0.5 million unexploited arbitrage opportunities on rollups. The Maximal Arbitrage Value (MAV) [21] on Arbitrum, Base, and Optimism pools varies from 0.03% to 0.05% of trading volume, whereas in ZKsync Era, it is approximately 0.25%. We show that episodes of price misalignment often last for 10–20 blocks in rollups; we further present a framework to capture only one MAV during each misalignment period, avoiding double-counting. The empirical MAV method that filters out repeated arbitrage results in 5×5$\times$5 × lower arbitrage profits compared to the Loss-Versus-Rebalancing (LVR) metric.
Cross-Rollup MEV. We estimate the potential for cross-rollup MEV by analyzing the price disparities among WETH-USDC Uniswap (v3) pools on major L2s. The arbitrageurs profits are influenced by the price impact (of AMM) and, thus, are the highest between the AMM pools with the highest token reserves.
2 Related Work
@@
There exists a substantial body of research on DeFi, particularly focused on AMM-DEXs. In particular, the systematization of knowledge (SoK) by Xu et al.  [46] and empirical investigations of Uniswap (v3) on the Ethereum blockchain [17, 20, 26, 27, 33] are prominent. Most of these studies evaluate the profitability of Liquidity Providers (LPs), specifically addressing impermanent loss. The analysis by Adams et al.  [1] analyzes the costs and dynamics of swaps on Uniswap (v3) on Ethereum. Our research extends this by breaking down swap fees on rollups and assessing the influence of blobs and block time on swap dynamics. Our findings are consistent with the simulated decomposition of swap fees in rollups presented in Gogol et al.  [22].
@@
The Loss Versus Rebalancing metric (LVR), which compares arbitrageur profits with those of LPs, was introduced by Milionis et al.  [37] and empirically measured by Fritsch et al.  [18]. LVR evaluates the LP portfolio against a rebalancing portfolio that perfectly hedges the market risk of the LP position. Our investigation quantifies the absolute profit of arbitrageurs, also known as the net LVR (per block), incorporating significant adjustments. First, LVR assumes that the price discrepancies between AMMs and target exchanges, such as CEXs, disappear after each block. Our findings confirm this assumption for Uniswap (v3) on Ethereum; however, the decay period of price discrepancies on rollups extends over multiple blocks. Our methodology adjusts the LVR to this scenario and prevents double counting of the same arbitrage opportunities. Second, LVR and the rebalancing portfolio assess the LP’s loss relative to arbitrageurs, thereby containing impermanent loss. Our metrics measure the absolute profit of the arbitrageur. Practically, arbitrageurs do not replicate the rebalancing portfolio but instead capitalize on arbitrage opportunities [24].
@@
The persistence of price disparities on AMMs within ZKsync Era, a ZK rollup, was initially reported by Gogol et al.  [21]. Our paper builds on that research by providing: i) a comprehensive analysis of rollups (Arbitrum, Base, Optimism, in addition to ZKsync Era) alongside Ethereum, ii) enhanced precision in market price accuracy and per-block analysis, improving from a 1-minute to a 1-second resolution, and iii) an in-depth examination of Uniswap (v3), focusing on the most actively traded pools.
@@
Heimbach et al.  [24] focus on the estimation of arbitrage executed through MEV auctions on Ethereum, specifically addressing non-atomic arbitrage opportunities between AMMs and CEXs. Our research extends this investigation by examining rollups, where MEV auctions have not yet been implemented. Similarly, we establish the minimum threshold for arbitrage opportunities and derive the arbitrage value using AMM formulas. However, our work seeks to quantify the unexploited arbitrage, which occasionally persisted for several minutes, in contrast to the arbitrage realized via MEV on Ethereum. Other studies on arbitrage [4, 5, 31] focus primarily on AMM-DEX on Ethereum and/or cyclic arbitrage opportunities.
@@
Torres et al.  [43] is the first effort to quantify MEV in rollups. Despite the absence of MEV auctions, this research examines the heuristic for back-running arbitrage MEV, specifically focusing on swap events occurring within the same transaction in the block, within the context of cylindrical arbitrage. Our objective is to assess the potential for non-atomic arbitrage, involving CEXs, or cross-rollup price discrepancies.
@@
Finally, Mamageishvili et al.  [34] introduce a time boost mechanism within rollup sequencers. This mechanism facilitates back-running MEV while preventing front-running MEV. In subsequent work [35], a shared sequencer was proposed, enabling cross-rollup MEV auctions, such as those for arbitrage exploitation. However, these studies lack empirical quantification of arbitrage opportunities within rollups, a gap that this work seeks to address.
@@ 3 Background
There are two main approaches — Layer-1 (L1) and Layer-2 (L2) scaling — to tackle blockchain scalability challenges and the associated increases in gas fees. L1 scaling involves creating a new blockchain with its own consensus mechanisms and a unique physical infrastructure to support the network. In contrast, L2 scaling uses a different method: it carries out computations off the main blockchain while recording their results on the underlying L1 chain. The main L2 scaling solutions include state (payment) channels, plasma, validium, and rollups. While plasma and state channels aim to move both data and computation off-chain, rollups and validium are non-custodial: they shift computation off-chain but keep compressed transaction data (rollups) or validity proofs (validiums) on the underlying L1 chain. As a result, rollups and validium require fewer trust assumptions compared to state channels, plasma chains, or L1 networks, with data security and integrity being ensured by the security of the underlying chain.
@@
Rollup. A rollup functions as a virtual blockchain, producing transaction and blocks and recording them on the main chain as compressed batches. However, it operates with the risk of adversarial actions such as stopping block production, generating invalid blocks, withholding data, or engaging in other malicious activities [7]. To ensure the correctness of the transaction state in these untrusted environments, rollups use one of two approaches: the optimistic approach and zero-knowledge (ZK) proofs. Figure 5, in §0.A, illustrates the rollup architecture, highlighting its principal components: sequencers and verifiers. Sequencers consolidate transactions to be rolled up onto the L1 chain, thereby achieving gas fee reductions. Verifiers, which are smart contracts on L1, validate the transaction block created by the sequencer, thus ensuring its correctness. Furthermore, ZK-rollups utilize provers responsible for generating ZK validity proofs [8].
@@
Soft and Hard Finality. Transaction finality refers to the point at which the state of a transaction in the blockchain ledger becomes permanent. Within the context of rollups, there are two types of transaction finality: hard finality, which is the permanent state of a transaction on the L1 chain, and soft finality, which is the irreversible state of a transaction on the L2 chain. [47]